OpenStudy (anonymous):

9.5-7.1+2.4x0.5-1.3=

OpenStudy (anonymous):

With the equation just like that? 2.3 (assuming you are multiplying 2.4 and 0.5)

OpenStudy (anonymous):

\[9.5 - 7.1 + (2.4) \times (0.5) - 1.3\] Is this your equation?

OpenStudy (anonymous):

I suggest you use MDAS (Multiplication, division, addition and subtraction

OpenStudy (anonymous):

how did you get that after I multiplied I assumed I had to add 7.1 +1.2 next?

OpenStudy (anonymous):

Once you have multiplied o divided all terms, add and subtract all terms starting from the left.

OpenStudy (anonymous):

9.5−7.1+(2.4)×(0.5)−1.3 Start by multiplying the 2.4 times 0.5, like @ettenelalovesnutella said, you should get the multiplication out of the way first.

OpenStudy (anonymous):

Thanks after I got 1.2 you simply go left to right? I thought I would have to add 7.1 then subtract 9.5 and then 1.3? 9.5-7.1+1.2-1.3?

OpenStudy (anonymous):

Yes. addition or subtraction. whichever goes first.

OpenStudy (anonymous):

really confused now .. if addition comes first 2.3 not the answer?

OpenStudy (anonymous):

We simply mean that once you have multiplied 2.4 and 0.5, you would then go back to the beginning and start again, only this time add or subtract. Therefore: 9.5-7.1+1.2-1.3= 2.4+1.2-1.3= 3.6-1.3= 2.3

OpenStudy (anonymous):

can you see why I am saying what I am if addition comes before subtraction?

OpenStudy (anonymous):

i feeeel dumb...lol

OpenStudy (anonymous):

Do you understand now? And it's okay. We all have these moments.

OpenStudy (anonymous):

Mhm. It's okay. It's just a math problem... ;D